Twenty years​ ago, 51​% of parents of children in high school felt it was a serious problem that high school students were not being taught enough math and science. A recent survey found that

252 of 750 parents of children in high school felt it was a serious problem that high school students were not being taught enough math and science. Do parents feel differently today than they did twenty years​ ago?

Use the alpha equals 0.1α=0.1 level of significance.

Because np 0 left parenthesis 1 minus p 0 right parenthesisnp01−p0equals=nothing

equals=

less than<

not equals≠

greater than>

​10, the sample size is

greater than

less than

​5% of the population​ size, and the sample

can be reasonably assumed to be random,

cannot be reasonably assumed to be random,

is given to not be random,

is given to be random,

the requirements for testing the hypothesis

are not

are

satisfied.

​(Round to one decimal place as​ needed.)
Answer #1

np(1 - p) = 750 * 0.51 * (1 - 0.51) = 187.4

Hence,

The correct answers in blanks will be:

187.4; >; Less than; Can be reasonably assumed to be normal; are